OpenStudy (anonymous):

I really need help with this, please help me. FIND THE VERTEX OF THE PARABOLA! Y=X^2-6X+7

hero (hero):

Hint: x = -b/(2a)

OpenStudy (anonymous):

thankyou, but im not sure where to go with that ^^

OpenStudy (tkhunny):

Other Hint: Complete the Square.

OpenStudy (anonymous):

totally lost :/

hero (hero):

The general form of a parabola is y = ax^2 + bx + c In this case we have y = x^2 -6x + 7 a = 1 b = -6 c = 7 To find x, input given values into vertex formula x = -b/(2a): x = -(-6)/2(1) = 6/2 = 3 The x coordinate of the vertex is 3 To find the y-coordinate, substitute x = 3 into the equation y = x^2 - 6x + 7: y = (3)^2 - 6(3) + 7 y = 9 - 18 + 7 y = 9 + 7 - 18 y = 16 - 18 y = -2 Therefore the y coordinate of the vertex is -2 Thus the vertex of the parabola is (x,y) = (3,-2)

OpenStudy (anonymous):

this was EXTREMELY helpful! thankyou :)